# Wide stable neural networks: Sample regularity, functional convergence and Bayesian inverse problems

### v1

@@ Abstract.
We study the large-width asymptotics of random fully connected neural networks with weights drawn from $\alpha$-stable distributions, a family of heavy-tailed distributions arising as the limiting distributions in the Gnedenko-Kolmogorov heavy-tailed central limit theorem. We show that in an arbitrary bounded Euclidean domain $\mathcal{U}$ with smooth boundary, the random field at the infinite-width limit, characterized in previous literature in terms of finite-dimensional distributions, has sample functions in the fractional Sobolev-Slobodeckij-type quasi-Banach function space $W^{s,p}(\mathcal{U})$ for integrability indices $p<\alpha$ and suitable smoothness indices $s$ depending on the activation function of the neural network, and establish the functional convergence of the processes in $\mathcal{P}(W^{s,p}(\mathcal{U}))$. This convergence result is leveraged in the study of functional posteriors for edge-preserving Bayesian inverse problems with stable neural network priors.
@@ 1. Introduction
The Gaussian process behaviour at the infinite-width limit has since been studied for a variety of neural network architectures, and different modes of convergence, in an expansive collection of papers by numerous authors. Motivations, apart from the context of Bayesian learning proposed by Neal, include the understanding of different initialization schemes for training, the formal study of gradient descent based training dynamics via the so-called neural tangent kernel at the infinite-width limit, and quantitative estimation of the convergence rates. As examples we mention the papers [31, 21, 29, 41, 18, 10, 2, 6, 13], although this list is by necessity very incomplete.
@@ 1.2. The model and previous literature
The scaling factor $H^{-1/\alpha}$ above is chosen in spirit of the heavy-tailed central limit theorem [15, §35, Theorem 2], as in it is in essence what one would expect to give rise to a scaling limit as $H\to\infty$ in case $\varphi$ were a bounded function.
@@ 1.3. Main results and discussion
    in $\mathcal{P}(W^{s,p}(\mathcal{U}))$ as $H\to\infty$.
@@
where $\boldsymbol{\varepsilon}_{M}$ is an independent $\mathbb{R}^{M}$-valued noise term (for example multivariate Gaussian). For given observations $\boldsymbol{u}$, the likelihood $\pi(\boldsymbol{u}\;\big{|}\;\cdot\,)$ is under suitable mild assumptions a bounded and continuous function on $W^{s,p}(\mathcal{U})$, and so the posterior with prior $\mathcal{L}(f^{H})$, $H\in\mathbb{N}\cup\{\infty\}$, can be characterized as
@@
in $\mathcal{P}(W^{s,p}(\mathcal{U}))$ as $H\to\infty$. We refer to Section 3 for details, and a posterior convergence result (Theorem 3.2) for more general forward operators.
@@
    The upper bound $\alpha$ for $p$ comes essentially from the application of Fubini’s theorem, and the fact that $\alpha$-stable random variables are $p$-integrable precisely for $p<\alpha$. We do not know whether this is the best possible integrability range for the $s$-smoothness for the sample functions of $f^{\infty}$ for given $s$. However, in general we do not expect the permissible range of $ps$ to extend beyond $\alpha\lambda$. For if this were the case for $(d,\alpha,\lambda)=(1,1,1)$, then the infinite-width limit of Cauchy neural networks on the real line would have Hölder-continuous sample paths (see Remark 1.3 (v) above), which empirically seems doubtful.
@@
    For $p<d/(d+s)$, it is possible to define Besov-type spaces such as $W^{s,p}(\mathcal{U})$ in a variety of ways, but the wide unity of different characterizations in the literature breaks down. For example, the Fourier-analytical definition of these spaces [39] will in this case contain tempered distributions that do not admit representations as functions. A more direct definition like that in Appendix A will result in a space of functions that does not embed in the space of tempered distributions, complicating the analysis of smoothness properties considerably [39, Section 2.2.3].
@@ 2. Sample regularity and functional convergence
    Part (ii) is of course obvious for $H<\infty$, in which case $f^{H}(x)$ is a continuous function of the input $x$ and a finite number of continuously distributed parameters.
@@
Let $\alpha$, $\beta$, $\lambda$ and $p$ be as in Proposition 2.3. Then for any measurable subset $A$ of $\mathcal{U}$ with nonzero measure, and any $n\in\mathbb{N}$, there exist points $(x^{n}_{j})_{j=1}^{\kappa(n)}\subset\mathcal{U}$ and scalars $(\theta^{n}_{j})_{j=1}^{\kappa(n)}\subset\mathbb{R}_{+}$ independent of $H$ such that
@@
Concerning part (ii), we recall from Proposition 2.6 that the local averages of $f^{H}$ convergence in distribution to those of $f^{\infty}$ as $H\to\infty$. By Proposition B.4 in Appendix B, the collection of local averages is a separating family for $\mathcal{P}(W^{s,p}(\mathcal{U}))$, so it follows that any $d$-convergent subsequence of $(\mathcal{L}(f^{H}))_{H\in\mathbb{N}}$ must converge to $\mathcal{L}(f^{\infty})$. This together with Proposition 2.5 then implies that the sequence $(\mathcal{L}(f^{H}))_{H\in\mathbb{N}}$ must itself converge to $\mathcal{L}(f^{\infty})$. ∎
@@ 3. Bayesian posterior analysis
Let $\varphi$ and $(p,s)$ be as in Theorem 1.1, and let $\boldsymbol{u}\in\mathbb{R}^{N}$. Assume that the density function $\rho_{M}$ of the noise term $\boldsymbol{\varepsilon}_{M}$ in (10) is bounded and Hölder-continuous with some exponent $\lambda_{\rho}\in(0,1]$, and that its support is the entirety of $\mathbb{R}^{M}$. Assume that the function $G\colon\mathbb{R}^{N}\times W^{s,p}(\mathcal{U})\to\mathbb{R}^{M}$ in (12) continuous, and uniformly Hölder-continuous with some exponent $\lambda_{G}\in(0,1]$ in the $\mathbb{R}^{N}$-component.
@@ 4. Deeper architectures
Let $\mathcal{U}$ and $\varphi$ be as in Theorem 1.1, with $\varphi$ Hölder-continuous (i.e. $\lambda=1$), and let $\alpha\in(d/(d+1),2)$. Then the statements of Theorems 1.1 and 3.2 continue to hold with $f^{\boldsymbol{H}}$ in place of $f^{H}$, and limiting process $f^{\infty}$ for $\min(H_{1},\cdots,H_{L})\longrightarrow\infty$ described by (18).

### v2

@@ Abstract.
We study the large-width asymptotics of random fully connected neural networks with weights drawn from $\alpha$-stable distributions, a family of heavy-tailed distributions arising as the limiting distributions in the Gnedenko-Kolmogorov heavy-tailed central limit theorem. We show that in an arbitrary bounded Euclidean domain $\mathcal{U}$ with smooth boundary, the random field at the infinite-width limit, characterized in previous literature in terms of finite-dimensional distributions, has sample functions in the fractional Sobolev-Slobodeckij-type quasi-Banach function space $W^{s,p}(\mathcal{U})$ for integrability indices $p<\alpha$ and suitable smoothness indices $s$ depending on the activation function of the neural network, and establish the functional convergence of the processes in the space of probability measures on $W^{s,p}(\mathcal{U})$. This convergence result is leveraged in the study of functional posteriors for edge-preserving Bayesian inverse problems with stable neural network priors.
@@ 1. Introduction
The Gaussian process behaviour at the infinite-width limit has since been studied for a variety of neural network architectures, and different modes of convergence, in an expansive collection of papers by numerous authors. Motivations, apart from the context of Bayesian learning proposed by Neal, include the understanding of different initialization schemes for training, the formal study of gradient descent based training dynamics via the so-called neural tangent kernel at the infinite-width limit, and quantitative estimation of the convergence rates. As examples we mention the papers [31, 21, 29, 41, 18, 10, 2, 6, 13], although this list is necessarily incomplete.
@@ 1.2. The model and previous literature
The scaling factor $H^{-1/\alpha}$ above is chosen in the spirit of the heavy-tailed central limit theorem [15, §35, Theorem 2], as in it is in essence what one would expect to give rise to a scaling limit as $H\to\infty$ in case $\varphi$ were a bounded function.
@@ 1.3. Main results and discussion
    in the space of Borel probability measures on $W^{s,p}(\mathcal{U})$ as $H\to\infty$.
@@
where $\boldsymbol{\varepsilon}_{M}$ is an independent $\mathbb{R}^{M}$-valued noise term (for example multivariate Gaussian). For given observations $\boldsymbol{u}$, the likelihood $\pi(\boldsymbol{u}\;\big{|}\;\cdot\,)$ is under suitable mild assumptions a bounded and continuous function on $W^{s,p}(\mathcal{U})$, and so the posterior induced by the prior $\mathcal{L}(f^{H})$, $H\in\mathbb{N}\cup\{\infty\}$, can be characterized as
@@
as $H\to\infty$. We refer to Section 3 for details, and a posterior convergence result (Theorem 3.2) for more general forward operators.
@@
    The upper bound $\alpha$ for $p$ comes essentially from the application of Fubini’s theorem, and the fact that $\alpha$-stable random variables are $p$-integrable precisely for $p<\alpha$. We do not know whether this is the best possible integrability range for the $s$-smoothness for the sample functions of $f^{\infty}$ for given $s$. However, in general we do not expect the permissible range of $ps$ to extend beyond $\alpha\lambda$. For if this were the case for $(d,\alpha,\lambda)=(1,1,1)$, then the infinite-width limit of Cauchy neural networks on the real line would have Hölder-continuous sample paths (see Remark 1.3 (v) above), which appears empirically unlikely.
@@
    For $p<d/(d+s)$, it is possible to define Besov-type spaces such as $W^{s,p}(\mathcal{U})$ in a variety of ways, but the wide unity of different characterizations in the literature breaks down. For example, the Fourier-analytical definition of these spaces [39] will in this case contain tempered distributions that do not admit representations as functions. A more direct definition, like the one preceding the statement of the Theorem, will result in a space of functions that does not embed in the space of tempered distributions, complicating the analysis of smoothness properties considerably [39, Section 2.2.3].
@@ 2. Sample regularity and functional convergence
    Part (ii) is, of course, obvious for $H<\infty$, in which case $f^{H}(x)$ is a continuous function of the input $x$ and a finite number of continuously distributed parameters.
@@
Let $\alpha$, $\beta$, $\lambda$ and $p$ be as in Proposition 2.3. Then for any measurable subset $A$ of $\mathcal{U}$ with nonzero measure, and any $n\in\mathbb{N}$, there exist points $(x^{n}_{j})_{j=1}^{\kappa(n)}\subset\mathcal{U}$ and scalars $(\theta^{n}_{j})_{j=1}^{\kappa(n)}\subset\mathbb{R}_{+}$ independent of $H\in\mathbb{N}\cup\{\infty\}$ such that
@@
Concerning part (ii), we recall from Proposition 2.6 that the local averages of $f^{H}$ convergence in distribution to those of $f^{\infty}$ as $H\to\infty$. By Proposition B.4 in Appendix B, the collection of local averages is a separating family for $\mathcal{P}(W^{s,p}(\mathcal{U}))$, so it follows that any $w$-convergent subsequence of $(\mathcal{L}(f^{H}))_{H\in\mathbb{N}}$ must converge to $\mathcal{L}(f^{\infty})$. This together with Proposition 2.5 then implies that the sequence $(\mathcal{L}(f^{H}))_{H\in\mathbb{N}}$ must itself converge to $\mathcal{L}(f^{\infty})$. ∎
@@ 3. Bayesian posterior analysis
Let $\varphi$ and $(p,s)$ be as in Theorem 1.1, and let $\boldsymbol{u}\in\mathbb{R}^{N}$. Assume that the density function $\rho_{M}$ of the noise term $\boldsymbol{\varepsilon}_{M}$ in (10) is bounded and Hölder-continuous with some exponent $\lambda_{\rho}\in(0,1]$, and that its support is the entirety of $\mathbb{R}^{M}$. Assume that the function $G\colon\mathbb{R}^{N}\times W^{s,p}(\mathcal{U})\to\mathbb{R}^{M}$ in (12) is continuous, and uniformly Hölder-continuous with some exponent $\lambda_{G}\in(0,1]$ in the $\mathbb{R}^{N}$-component.
@@ 4. Deeper architectures
Let $\mathcal{U}$ and $\varphi$ be as in Theorem 1.1, with $\varphi$ Lipschitz-continuous (i.e. $\lambda=1$), and let $\alpha\in(d/(d+1),2)$. Then the statements of Theorems 1.1 and 3.2 continue to hold with $f^{\boldsymbol{H}}$ in place of $f^{H}$, and limiting process $f^{\infty}$ for $\min(H_{1},\cdots,H_{L})\longrightarrow\infty$ described by (18).